OpenStudy (anonymous):

Please help confirm my answer: log_3(7x-1)-log_3(x+1)=log_3 5

OpenStudy (anonymous):

@Hero I have another one to confirm if you dont mind. (7x-1)/(x+1)=5 6x-1=5x+5 x=6

OpenStudy (anonymous):

@jim_thompson5910 can you confirm the answer for me

jimthompson5910 (jim_thompson5910):

no, 6 is not the answer

OpenStudy (anonymous):

yikes can you show me how to get to the answer

jimthompson5910 (jim_thompson5910):

log_3(7x-1)-log_3(x+1)=log_3 5 log_3[(7x-1)/(x+1) ] =log_3 5 (7x-1)/(x+1) = 5 7x-1 = 5(x+1) ... this is where you went wrong 7x - 1 = 5x + 5 7x - 5x = 5 + 1 2x = 6 x = 6/2 x = 3

jimthompson5910 (jim_thompson5910):

you somehow said 6x - 1 instead of 7x - 1